Futuristic was a newly setup firm in 2019. It is in the business of providing artwork. However, being in the business of standard art objects (non-essential items), it did not do well once it was hit by COVID-19. It has a selling price of ₹2,500 per piece and a variable cost of ₹1,000 per piece. It incurs an annual fixed cost of ₹30,00,000.

 

a. The manager of the business wishes to know-

  • the number of art pieces they must sell to be at the break-even point
  • the contribution margin
  • whether the firm is earning a profit or incurring a loss by selling 2,100 art pieces.

b. To combat the current difficult situation, the manager plans to curb the variable expenses and bring them to ₹500 per piece. Compute the new break-even point and contribution margin. Analyze and explain the movement in contribution margin to the manager.


1
Expert's answer
2021-11-01T17:34:02-0400

The break-even point is:

"BEP = 3,000,000\/(2,500 - 1,000) = 2,000" units.

The contribution margin is:

"CM = (2,500 - 1,000)\/2,500 = 0.6."